# COSINE-100 Collaboration

### v2

@@ III.2 Atomic ionization factor
The ionization factor describes the electronic structure of the target material, and different models can be considered depending on the deposited energies in the detector. The first studies conducted in xenon and argon detectors used to assume an analytic non-relativistic approach when computing $f_{ion}$, considering Slater-type orbitals for the bound electrons wavefunctions Agnes et al. (2018); Essig et al. (2012). This procedure leads to reasonable approximations when analyzing very low energy events (typically of O(1-10 eV)), in which few electrons are ionized. However, it neglects important phenomena as the non-point like behavior of nuclear charge for the electron wavefunction near the nucleus, and relativistic effects, which are important for higher energy analysis. For this situation, a relativistic description of the electronic structure is important for an accurate analysis.
@@
In this work, since the ROI is in the keV region, due to the 0.7 keV threshold of COSINE-100 crystals, sodium and iodine atoms can be considered to be free, as the typical crystal band structure energies are much lower than this Derenzo et al. (2017); Griffin et al. (2021); Roberts and Flambaum (2019). Therefore, including the ionization factor contribution of NaI(Tl) crystal structure would marginally impact the results of this work. In general, when calculating $f_{ion}$ for each atom isolated, it can be determined from the electron wavefunctions of each shell. However, very distinct results are obtained with different modeling of this atomic electronic structure. More details about the analytic models and a discussion and comparison between non-relativistic and relativistic methods of $f_{ion}$ computation are presented in Appendix A.
@@
$f_{ion}$ is dependent on both $E$ and $q$. Typically, for energy depositions in the keV region, there is an important contribution of $f_{ion}$ in the range of $q\gtrsim 1$ MeV to the event rate. For such high momentum transfer, there is a high correlation of $f_{ion}$ with the radial electron wavefunction near the nucleus, where relativistic effects and correct nucleus size modeling are crucial. Hence, this analytical approach reproduces with good accuracy $f_{ion}$ only for small $q$, which is suitable for very low energy deposition in the 1-100 eV range.
@@
A more complete model for a keV region analysis should account for relativistic effects. A possible procedure is to numerically solve the Dirac equation, correctly considering the atomic potential for different distances of the electron from the nucleus by using the relativistic Hartree-Fock approximation Roberts et al. (2016). This approach has been used in the DAMA/LIBRA results interpretation under the WIMP-electron scattering model. Computation of $f_{ion}$ for sodium and iodine atoms, considering them as free, was performed by authors of the DAMA/LIBRA results interpretation study, and are publicly available in Roberts and Caddell (2023). The electronic structure and methods for calculations are thoroughly discussed in Roberts et al. (2016); Roberts and Flambaum (2019). These results are used for the WIMP-electron interaction spectra generation in the analysis presented in this work.
A comparison between the expected spectra in NaI using the analytic non-relativistic approach, described in the Appendix, and the relativistic method for obtaining $f_{ion}$ is presented in Fig. 3. In both cases, the iodine contribution is dominant compared to sodium, even though both are considered in this analysis.
The expected event rates below 0.4 keV are compatible for both $f_{ion}$ models, since only low and intermediate $q$ values (typically below 1 MeV) are dominant in this energy region. However, in the most relevant region for this analysis, above the 0.7 keV threshold, the disagreement is around one order of magnitude. Consequently, the analytic procedure underestimates the number of events, reducing the analysis sensitivity. Another notable behavior is the slight shift in ionization energies between both models. Due to the relativistic effects, electron binding energies undergo a small correction in calculations for the non-relativistic model.
@@
Figure 3: Computed spectra in NaI considering the non-relativistic and relativistic methods. In the blue curves, $m_{\chi}=1$ GeV is considered, and in the orange curves, $m_{\chi}=10$ GeV is considered. Solid lines represent the computation of spectra with the relativistic approach for $f_{ion}$ calculation, while dashed lines represent spectra with the non-relativistic approach for $f_{ion}$ calculation.
@@ IV.1 Event selection
With the objective of greatly reducing the number of noise events in the ROI, a deep learning method based on the multilayer perceptron (MLP) neural network was used, allowing the reduction of analysis threshold from the previous 1 keV to 0.7 keV, maintaining an excellent event selection efficiency at the threshold level. Pulse shape discrimination likelihood parameters were introduced as inputs to MLP. The waveforms mean time and amplitudes were evaluated to obtain a PSD parameter. The fast Fourier transform of waveforms has also been included in the generation of another PSD parameter based on the waveforms power spectra. Scintillation events from calibration using the 0.87 keV X-ray emission from the electron capture process in a 22Na source, and noise events from PMT origin were inputs to the MLP in the sample training. A MLP score based on the likelihood parameters cited was calculated for each signal, and the event selection cut determined following the criteria that less than 1% of the selected events were noise signals Yu et al. (2024). The efficiency of this noise and scintillation event selection is presented and discussed in the most recent COSINE-100 WIMP-nucleon interaction search Yu et al. .
@@
IV.2 COSINE-100 crystals response and WIMP-electron scattering spectra
In order to generate the expected spectra in COSINE-100 crystals for the WIMP-electron scattering model, it is essential to properly consider the detector’s energy resolution and the non-linear proportionality between the measured charge in the PMTs and deposited energy at low photoelectrons level. As a consequence of the small energy deposition of WIMPs on electrons, the number of events should be higher below the 8 photoelectrons threshold. The detector performance in this region has to be rigorously studied, since these events may induce signals exceeding the threshold due to the energy resolution effect.
The crystals resolution and non-proportionality at very low energies were carefully examined by data from calibrations using the 0.87 keV X-ray emission of the 22Na source and background signals from the 3.2 keV X-ray peak of internal 40K Lee et al. (2024). The detector behavior has been reproduced down to one photoelectron level in the PMTs waveform simulation Choi et al. (2024a). Accordingly to these studies, the following scaled Poisson distribution describes the crystals energy resolution for few photoelectrons yield:
@@ IV.2 COSINE-100 crystals response and WIMP-electron scattering spectra
where $q_{raw}$ is the deposited charge in the detector, $q_{vis}$ is the visible charge, and the $s$ parameter is determined from calibration Lee et al. (2024).
The generated WIMP-electron spectra in COSINE-100 NaI(Tl) crystals, corrected by event selection efficiency, detectors’ non-proportionality and energy resolution are presented in Fig. 4, assuming the heavy mediator scenario. Most signals are expected to be detected with energies near the 0.7 keV threshold for all WIMP masses, which makes the ROI of this analysis to be at very low energy depositions. Hence, we considered only data up to 100 photoelectrons (equivalent to about 8 keV) in the data fit. Also, the expected number of events is highest for $m_{\chi}\approx$ 0.6 GeV when considering the heavy mediator case, and for $m_{\chi}\approx$ 0.2 GeV when considering the light mediator case. Therefore, the COSINE-100 sensitivity is expected to be higher near these WIMP masses. Additionally, it is important to note that the first energy bin, for lower mass WIMPs and especially for the light mediator scenario, has multiple times the number of events of the other bins of the expected signal, making the understanding of background near the threshold very important to enhance the analysis sensitivity for WIMP masses in the MeV range.
Figure 4: Expected event rate in COSINE-100 NaI(Tl) detectors considering a heavy mediator and 5 different WIMP masses. The curves represent an average over the 5 analyzed crystals. $\overline{\sigma}_{e}$ is fixed at $10^{-38}$ cm2.
@@ IV.3 Background components
[…]
Systematic uncertainties presented in the background model have different sources, with the most relevant at the ROI being uncertainties in the: MLP event selection efficiencies; determination of the precise location of background sources in the crystals’ PMTs; evaluation of the parameters that describe the crystals’ energy resolution distribution of Equation 6; non-proportionality impact on the calibration; measurement of the activity of each background component; and depth of the surface 210Pb in the crystals. The background model used in this analysis is the same described in the recent COSINE-100 WIMP search Yu et al. , with the only difference being that a scaled Poisson distribution is considered instead of a Gaussian distribution when correcting the spectra of background components by the energy resolution effect. This is important to match the energy resolution correction applied to the WIMP-electron interaction signals with the correction of background components spectra. The main different results to the final background model are generated by radioisotopes that produce few photoelectrons, since the shape of the scaled Poisson function differs from the Gaussian function at very low energy depositions. Thus, 22Na, 109Cd, and 113Sn cosmogenic isotopes are the components which differed the most with the modification of energy resolution function, due to their emission of X-rays and Auger electrons near the 0.7 keV threshold. Nevertheless, when compared to the official COSINE-100 crystals background model Yu et al. (2025), the change in the energy resolution distribution marginally increased the number of events of the background model spectrum, up to a maximum of 2% in the first energy bin.
IV.4 Spectral fit
A Bayesian analysis approach was used in the spectra fit to the 2.82 years data-set process. The posterior distribution for the WIMP-electron signal was obtained from a likelihood function based on Poissonian probabilities, determined by a Markov Chain Monte Carlo (MCMC) based algorithm. This analysis method is the same adopted in other DM searches in COSINE-100 Adhikari et al. (2023a, b, c, 2020), and is thoroughly described in Ref. Adhikari et al. (2021b).
The fit was performed for several WIMP masses, ranging from 30 MeV to 5 GeV. The mean background activities and statistical uncertainties, determined from each crystal background model Yu et al. (2025), were used to produce the Gaussian priors. Systematic uncertainties were included as nuisance parameters also with Gaussian priors. Each crystal background is allowed to independently fluctuate in the simultaneous fit performed, but the $\overline{\sigma}_{e}$ distribution is marginalized over all 5 crystals. No evidence of excess events consistent with expected spectra from both the light and heavy mediator models was found. The resulting fit for a 0.6 GeV WIMP and assuming a heavy mediator is presented in Fig. 6, along with the expected observed data for $\overline{\sigma}_{e}=9.2\times 10^{-38}$ cm2.
[…]
After marginalization, the resulting posterior probability distribution function (PDF) for the WIMP-electron signal strength is used to set a 90% confidence level (CL) upper limits on $\overline{\sigma}_{e}$. Fig. 7 shows the posterior distribution for the 0.6 GeV WIMP data fit.
@@ IV.4 Spectral fit
In the light mediator scenario, the distance between the set upper limits and the median sensitivity remains almost constant for any WIMP mass. This effect reflects the small change in the spectra shape for low WIMP masses, and particularly for the light mediator model, a spectrum shape variation is only relevant above 10 GeV.
The upper limits set from this COSINE-100 data analysis for a light and heavy mediator scenarios do not explore regions unconstrained by the most sensitive experiments for the WIMP-electron interaction model, due to their capability of reaching very low threshold (down to a single electron ionization Aprile et al. (2022)). However, the reported COSINE-100 constraints completely exclude the best fit regions for the DAMA/LIBRA modulation results assuming any vector mediator mass, and set the current most stringent upper limits for the WIMP-electron scattering model for a NaI(Tl) experiment.

### v3

@@ III.2 Atomic ionization factor
Different descriptions are considered when determining the $f_{ion}$ parameter in light DM-electron interaction studies, and an accurate description is essential depending on the target and energies under study, as very divergent results can be obtained [47].
@@
The ionization factor describes the electronic structure of the target material, and different models can be considered depending on the deposited energies in the detector. The first studies conducted in xenon and argon detectors used to assume an analytic non-relativistic approach when computing $f_{ion}$, considering Slater-type orbitals for the bound electrons wavefunctions [48, 20]. This procedure leads to reasonable approximations when analyzing very low energy events (typically of O(1-10 eV)), in which few electrons are ionized. However, it neglects important phenomena as the non-point like behavior of nuclear charge for the electron wavefunction near the nucleus, and relativistic effects, which are important for higher energy analyses. For the case of this work, since the ROI is in the keV region due to the 0.7 keV analysis threshold of COSINE-100 crystals, a relativistic description of the electronic structure is important for an accurate analysis of the DM-electron scattering [31].
@@
For this analysis, sodium and iodine atoms can be considered to be free, as the typical crystal band structure energies are much lower than the keV scale [49, 50, 32]. Therefore, including the ionization factor contribution of NaI(Tl) crystal structure would marginally impact the results of this work. In general, when calculating $f_{ion}$ for each atom isolated, it can be determined from the electron wavefunctions of each shell. However, very distinct results are obtained with different modeling of this atomic electronic structure. More details about the analytic models and a discussion and comparison between non-relativistic and relativistic methods of $f_{ion}$ computation are presented in Appendix A.
@@
$f_{ion}$ is dependent on both $E$ and $q$. Typically, for energy depositions in the keV region, there is an important contribution of $f_{ion}$ in the range of $q\gtrsim 1$ MeV to the event rate. For such high momentum transfer, there is a high correlation of $f_{ion}$ with the radial electron wavefunction near the nucleus, where relativistic effects and correct nucleus size modeling are crucial. Hence, this analytical approach reproduces $f_{ion}$ with good accuracy only at small $q$, suitable for very low energy depositions in the 1–100 eV range in materials whose atoms interaction is negligible, such as noble-gas targets. This approximation, however, is not valid for crystalline materials such as Ge, Si, or NaI(Tl).
@@
A more complete model for a keV region analysis should account for relativistic effects. A possible procedure is to numerically solve the Dirac equation, correctly considering the atomic potential for different distances of the electron from the nucleus by using the relativistic Hartree-Fock approximation [31]. This approach has been used in the DAMA/LIBRA results interpretation under the DM-electron scattering model. Computation of $f_{ion}$ for sodium and iodine atoms, considering them as free, was performed by authors of the DAMA/LIBRA results interpretation study, and are publicly available in [51]. The electronic structure and methods for calculations are thoroughly discussed in [31, 32]. These results are used for the DM-electron interaction spectra generation in the analysis presented in this work.
A comparison of the expected spectra in NaI obtained using the analytic non-relativistic approach, described in Appendix A, and the relativistic method for obtaining $f_{ion}$ is shown in Fig. 3. A notable difference between both approaches is a slight shift in the ionization energies due to distinct treatments of electronic wave functions. Relativistic effects lead to small corrections to the electron binding energies relative to calculations for the non-relativistic model, as can be seen by comparing the electron binding and ionization energies reported in Refs. [52, 31]. In Fig. 3, however, the ionization energies are fixed to the values of the relativistic approach to enable a more direct comparison between the two models.
Regardless of the chosen treatment of atomic electrons, the iodine contribution is dominant compared to sodium for the ROI of this analysis. At 0.1 keV, sodium reaches its maximum contribution of 12% to the total event rate, and decreases down to 0.3% at 0.7 keV. Due to convolution of the raw spectrum of Fig. 3 with the detector energy resolution function, a small part of events below the 0.7 keV threshold can be detected. Hence, both iodine and sodium atomic ionization factors have been considered in this analysis.
Thallium atoms in NaI(Tl) crystals can also be ionized by the DM particle and contribute to the expected event rate. However, even though thallium has roughly 53% more electrons than iodine atoms, its concentration in COSINE-100 NaI(Tl) crystals is approximately 0.1 mol% [53, 54], which makes its contribution to be insignificant in this analysis. Naive calculations assuming the non-relativistic approach have shown that the maximum thallium contribution is 2% at 0.9 keV, and therefore it was neglected.
@@
Figure 3: Computed spectra in NaI using the analytic non-relativistic and relativistic methods, considering a heavy mediator. In the blue curves, $m_{\chi}=1$ GeV is considered, and in the orange curves, $m_{\chi}=10$ GeV is considered. Solid lines represent the computation of spectra with the relativistic approach for $f_{ion}$ calculation, while dashed lines represent spectra with the non-relativistic approach for $f_{ion}$ calculation. The kinks observed between 0.05 keV and 0.3 keV and between 0.8 keV and 1.2 keV arise from electron ionization energies associated with different shells of Na and I atoms.
For lower energy depositions, the raw event rates using both $f_{ion}$ models are closer to each other, since only low and intermediate $q$ values (typically below 1 MeV) are dominant at this energy region. It is important to note that at energy depositions typically below 0.1 keV, the crystal band structure of NaI(Tl) becomes relevant, and the treatment of isolated sodium and iodine atoms adopted here cannot be used [49]. In the most relevant energy region for this analysis, near the 0.7 keV threshold, the disagreement reaches approximately one order of magnitude. As a consequence, the analytic approach underestimates the expected event rate, leading to a reduction in sensitivity. Hence, the use of the more accurate relativistic method is essential for describing energy depositions in the keV region.
@@ IV Data analysis
To search for DM–electron scattering events in the COSINE-100 data, a detailed treatment of the low energy signals, along with the understanding of the crystals’ response and background components in the ROI is required.
The following section IV.1 describes the selection of low energy COSINE-100 data suitable for the search for the DM–electron interaction. Section IV.2 discusses the expected response of the NaI(Tl) crystals to DM electron signals and its impact on this analysis. Section IV.3 presents the background model, including a discussion of the main components and uncertainties relevant to the ROI. Section IV.4 describes the fit of the DM–electron energy spectra to the COSINE-100 data.
@@ IV.1 Event selection
With the objective of greatly reducing the number of noise events in the ROI, a deep learning method based on the multilayer perceptron (MLP) neural network was used [55], allowing the reduction of analysis threshold from the previous 1 keV to 0.7 keV, maintaining an excellent event selection efficiency at the threshold level. Pulse shape discrimination likelihood parameters were introduced as inputs to MLP. The waveforms mean time and amplitudes were evaluated to obtain a PSD parameter. The fast Fourier transform of waveforms has also been included in the generation of another PSD parameter based on the waveforms power spectra. Scintillation events from a one month calibration campaign using an external 22Na source, together with PMT-induced noise events, were used as inputs for the MLP sample training. Pure scintillation signals were selected by requiring multiple-hit events in the crystals. The 22Na decay produces a 1274 keV gamma ray from the de-excitation of the 22Ne daughter nucleus, as well as two 511 keV gamma rays from positron–electron annihilation [56]. Therefore, pure low-energy scintillation signals from Compton scattering in a given crystal could be selected by requiring a coincidence with high-energy events in another crystal. A MLP score based on the likelihood parameters was calculated for each signal, and the event selection cut determined following the criteria that less than 1% of the selected events were noise signals. The efficiency of this noise and scintillation event selection is discussed in details in the most recent COSINE-100 WIMP-nucleon interaction search [35] and in the study of software techniques for lowering COSINE-100 detector threshold [33]. In addition, the trigger efficiency is verified using a waveform simulation of the PMTs signals [57] and is found to be consistent with unity at the current 8 photoelectrons threshold.
@@
Figure 4: Event selection efficiencies as a function of energy for the 5 crystals used in this analysis. The gray shaded region indicates the energy range below the 8 photoelectron analysis threshold. The dashed black line denotes 100% efficiency.
IV.2 COSINE-100 crystals response and DM-electron scattering spectra
In order to generate the expected spectra in COSINE-100 crystals for the DM-electron scattering model, it is essential to properly consider the detector’s energy resolution and the non-linear proportionality between the measured charge in the PMTs and deposited energy at low photoelectrons level. As a consequence of the small energy deposition of DM on electrons, the number of events should be higher below the 8 photoelectrons threshold. The detector performance in this region has to be rigorously studied, since these events may induce signals exceeding the threshold due to the energy resolution effect.
The crystals resolution and non-proportionality at very low energies were carefully examined by data from calibrations using the 0.87 keV X-ray emission of the 22Na isotope and the 3.2 keV X-ray peak of the 40K isotope, both present as internal contaminants of the crystals [58]. The detector behavior has been reproduced down to one photoelectron level in the PMTs waveform simulation [57]. Accordingly to these studies, the following scaled Poisson distribution describes the crystals energy resolution for few photoelectrons yield:
@@ IV.2 COSINE-100 crystals response and DM-electron scattering spectra
where $q_{raw}$ is the deposited charge in the detector, $q_{vis}$ is the visible charge, and the $s$ parameter is determined from calibration [58].
The generated visible spectra in COSINE-100 NaI(Tl) crystals, corrected by event selection efficiency, detectors’ non-proportionality and energy resolution are presented in Fig. 5, assuming the heavy mediator scenario. Most signals are expected to be detected with energies near the 0.7 keV threshold for all DM masses, which makes the ROI of this analysis to be at very low energy depositions. Hence, we considered only data up to 100 photoelectrons (equivalent to approximately 7.8 keV) in the data fit. Also, due to the discussed detector response, the expected number of events is highest for $m_{\chi}\approx$ 0.6 GeV when considering the heavy mediator case, and for $m_{\chi}\approx$ 0.2 GeV when considering the light mediator case. Therefore, the COSINE-100 sensitivity is expected to be higher near these DM masses. In addition, for lower DM masses, the expected event rate in the first energy bin is several times larger than that in the higher energy bins. As can be observed in Fig. 3, light DM particles in the halo do not have sufficient kinetic energy to deposit keV scale energies in the detector, resulting in most events occurring below the COSINE-100 threshold. Nevertheless, due to the energy resolution of the detector, a fraction of these events can be reconstructed above 0.7 keV, while retaining a quickly falling spectral shape. Consequently, an accurate understanding of the background near threshold is crucial for enhancing the sensitivity of the analysis to DM particles in the MeV mass range.
Figure 5: Expected event rate in COSINE-100 NaI(Tl) detectors considering a heavy mediator and 5 different DM masses. The curves represent an average over the 5 analyzed crystals. $\overline{\sigma}_{e}$ is fixed at $10^{-38}$ cm2.
@@ IV.3 Background components
The activities of each component that constitute the COSINE-100 crystals background are known from a detailed detector simulation based on Geant-4 [59]. The detector exposure time of the data-set used for the background modeling is identical to that considered in this analysis. The activity of each isotope is determined through a combined fit to single-hit events in the energy range from 6 keV to 4000 keV and to multiple-hit events from 0.7 keV to 4000 keV. Single-hit events in the [0.7–6] keV region are excluded from the background Monte Carlo (MC) fit in order to preserve blindness to potential low energy DM signals, such as the expected from DM–electron scattering [60].
@@
Table 1: Contribution of each systematic error component of the background model in the ROI
Systematics component Contribution ratio (%)
[0.7 - 1.5] keV [1.5 - 4] keV [4 - 7.8] keV
Event selection efficiency 67.59 12.52 11.71
PMT background sources location 4.03 8.09 21.31
Energy resolution parameters 0.23 2.48 6.34
Activity of background components 2.18 11.57 28.23
Depth of surface 210Pb 25.98 65.34 32.40
The background model used in this analysis is the same described in the recent COSINE-100 WIMP search [35], with the only difference being that a scaled Poisson distribution is considered instead of a Gaussian distribution when correcting the spectra of background components by the energy resolution effect. This is important to match the energy resolution correction applied to the DM-electron interaction signals with the correction of background components spectra. The main different results to the final background model are generated by radioisotopes that generate few photoelectrons, since the shape of the scaled Poisson function differs from the Gaussian function at very low energy depositions. Thus, 22Na, 109Cd, and 113Sn cosmogenic isotopes are the components which differed the most with the modification of the energy resolution function, due to their emission of X-rays and Auger electrons near the 0.7 keV threshold. Nevertheless, when compared to the official COSINE-100 crystals background model [60], the change in the energy resolution distribution marginally increased the number of events in the background model spectrum, up to a maximum of 2% in the first energy bin.
IV.4 Spectral fit
A Bayesian analysis approach was used in the spectra fit to the 2.82 years data-set process. The posterior distribution for the DM-electron signal was obtained from a likelihood function based on Poisson probabilities, determined by a Markov Chain Monte Carlo (MCMC) based algorithm. This analysis method is the same adopted in other DM searches in COSINE-100 [61, 62, 63, 64], and is thoroughly described in Ref. [65].
The fit was performed for several DM masses, ranging from 30 MeV to 5 GeV. The mean background activities and statistical uncertainties, determined from each crystal background model [60], were used to produce the Gaussian priors. Systematic uncertainties were included as nuisance parameters also with Gaussian priors. Each crystal background is allowed to independently fluctuate in the simultaneous fit performed, but the $\overline{\sigma}_{e}$ distribution is marginalized over all 5 crystals. No evidence of excess events consistent with expected spectra from both the light and heavy mediator models was found. The resulting fit for a 0.6 GeV DM and assuming a heavy mediator is presented in Fig. 7, along with the expected observed data for $\overline{\sigma}_{e}=9.2\times 10^{-38}$ cm2.
[…]
After marginalization, the resulting posterior probability distribution function (PDF) for the DM-electron signal strength is used to set a 90% confidence level (CL) upper limits on $\overline{\sigma}_{e}$. Fig. 8 shows the posterior distribution for the 0.6 GeV DM data fit.
[…]
For lighter DM masses, in both light and heavy mediator scenarios, the distance between the set upper limits and the median sensitivity remains almost constant for any DM mass. This effect reflects the small change in the spectra shape for low mass DM particles.
The upper limits set from this COSINE-100 data analysis for a light and heavy mediator scenarios do not explore regions unconstrained by the most sensitive experiments for the DM-electron interaction model, due to their capability of reaching very low threshold (down to a single electron ionization [19]). However, the reported COSINE-100 constraints completely exclude the best fit regions for DAMA/LIBRA modulation results assuming any vector mediator mass, and set the current most stringent upper limits for the DM-electron scattering model for a NaI(Tl) experiment.